OpenStudy (anonymous):

Rewrite with rational exponents. ^3sqrt xy^2 z

OpenStudy (anonymous):

\[^3\sqrt xy^2 z\] The square root goes all the way across

OpenStudy (anonymous):

\[\sqrt[3]{xy^2z}=(xy^2z)^{\frac{1}{3}}\]

OpenStudy (anonymous):

btw that is "cube root" not "square root"

OpenStudy (anonymous):

oops. Thanks for correcting me :)

OpenStudy (anonymous):

\[(^{7}\sqrt{2x ^{4}}y)^{11}\]